OpenStudy (anonymous):

given the trinomial 3x^2-6x+5, what is the value of the discriminant?

OpenStudy (campbell_st):

the discriminant comes from the general quadratic formula it is \[\Delta = b^2 -4ac\] so looking at the coefficients in your question a = 3, b = -6 and c = 5 just substitute and evaluate

OpenStudy (anonymous):

no i wrote the problem exactly like it said. the answer choices were: A= -66, B= -24, C= 8, D= 96 @campbell_st

OpenStudy (campbell_st):

so what answer did you get when you substituted \[\Delta = (-6)^2 - 4 \times 3 \times 5\]

OpenStudy (anonymous):

-600? @campbell_st

OpenStudy (campbell_st):

not quite do you know what \[(-6)^2 = ?\] it can be written as -6 x -6 =

OpenStudy (anonymous):

yeah its 36

OpenStudy (campbell_st):

so what is 36 - (4 * 3 * 5) =

OpenStudy (anonymous):

oh ok its -24. your awesome:)

OpenStudy (campbell_st):

glad to help...